OpenStudy (anonymous):

find the zeros of the function algebraically. f(x)= 2x^2-9/(3-x)

OpenStudy (anonymous):

f(0)=x

OpenStudy (anonymous):

(2x+3)(2x-3)/(3-x) = 0 (2x+3)(2x-3) = 0 2x+3=0 2x=-3 x=-3/2 similarly 2x=3 x=3/2 zerios are 3/2 and -3/2